ACM Email Room Staff Instructions

1. No matter how boneheaded someone seems, be polite to them.

2. Each time you enter the room, and at least once every 15 minutes, count the number of radio laptops and check that both hubs, the printer and the cables are all in place. If anything is missing:

(a) Do nothing that could put you in danger.

(b) Take a quick look around to see if you can find the missing item.

(c) Quickly call Susan Austic at the registration desk, explain what’s happened, and ask for reinforcements. The number is 227-0628.

(d) Close the door to the room. Do not admit anyone other than ACM or Hotel staff or the police. Ask people not to leave (but do not hinder them if they insist). If someone insists on leaving, get a good look at his or her face, write the name on his or her badge down, and see of he or she is carrying anything that could conceal the missing item.

(e) Call Joe Harrington at 5-5913 or his cell phone at 280-5990.

(f) Notify Peter Meixell at 4-2573.

(g) Notify Jeff Hetzel at 5-6383.

3. The room must be staffed at all times, without even one minute unstaffed. If you need to leave the room, call for a relief worker. Any ACM staffer will do. You can call the AV rooms, the registration desk, or Joe’s cell phone. Wait for them to arrive and count the laptops before you go and when you return. The same goes for shift changes, so please arrive a few minutes early for your shift. If you have a dire emergency, must leave, and cannot find a relief worker, ask everyone in the room to leave, count the laptops, and lock the door behind you.

4. When someone asks for an IP address, use the next available address on the ACM IP Address Log. Record the person’s machine type, name, and affiliation on the log. Select the appropriate Participant Setup sheet, write the IP number on it yourself, and give it to the person.

The last 16 IP addresses are temporary. If we have to use them, tell people that they are only good for one session and that others will use those addresses at other times, so they need to get their number reassigned each time they plug into the net.

5. Participants are responsible for setting up their own laptops. Don’t touch their machines. If someone with a personal laptop asks you how to set it up, ask them first whether they have been assigned an IP address by an email room staffer, and assign one if necessary. If they’ve been assigned a number and can’t get it to work by following the instructions, check that the cable is plugged into their machine and that the connector isn’t damaged. If there’s still a problem, you may talk them through the setup. Remember to make them record their original numbers. If it doesn’t work, however, it’s not your responsibility to make it work.

6. Nobody but ACM or Hotel staff loads the printer. If you see someone messing with it, ask them what they need and do it for them if it’s reasonable.

7. Transparencies are only for presentations to be made at ACM. Hand out only the number requested and no more than 20 to a person. Load them into the printer yourself.

8. Every 20 minutes write down how many people are:

• using one of our machines

• using machines of their own

• waiting to use one of our machines

• waiting for a cable to free up